# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v2

@@ 3.7 Baseline models
In all our experiments, we will compare our MapFormer models to a classic Transformer architecture with static positional encoding, as well as two architectures learning input-dependent positional encoding serving as baselines for structure learning models.
Rotary Position Embedding (RoPE)
@@
is used as our baseline Transformer since it is the most popular (fixed) position encoding architecture, and is analogous to our MapWM model without input selectivity, theoretically incapable of learning input-dependent matrices and structure-based problems su2023roformerenhancedtransformerrotary.
Contextualized Positional Encoding (CoPE)
is an enhanced Transformer architecture that updates its relative positions between queries and keys with a cumulative gating, akin to Mamba2 dao2024transformersssmsgeneralizedmodels. This allows CoPE to ignore irrelevant tokens and perfectly generalize OOD to several structure-dependent tasks. The cumulative gating of CoPE reminds path integration, but since the sigmoid gate can only take values between 0 and 1, the model cannot represent an action and its inverse, and should not be able to represent trajectories where the agent moves both forward and backward. Moreover, since CoPE only compresses relative distances, it should not be able to represent structure in higher dimensions, where temporal order does not matter golovneva2024contextualpositionencodinglearning.
Contextualized Equivariant Position Embedding (TAPE)
@@
learns to modify its positional encoding after the attention mechanism, with input-dependent MLPs, resembling the input-dependent matrices of EM/WM actionable models Whittington2025. However, TAPE exhibits several issues: First, updating structure after the attention means that the latter cannot benefit from input-dependent positional encoding in the first layer, meaning that structure-based recall with a single-layer model is impossible. Second, updating values before updating structure means that values could be "polluted" by structurally irrelevant tokens, adding unwanted noise to the data. Third, TAPE’s positional embedding are updated based on the attention scores and not path integration, giving no theoretical guaranty to preserve structure over long ranges zhu2025rethinkingaddressinglanguagemodels.
4 Implementation details
Rotations can be ensured by enforcing $S:=\overline{S}-\overline{S}^{\top}$ to be skew-symmetric, but since a single generator matrix $S$ can only generate rotations along a single axis (see sec.10.4.2 for further details), choosing blocks of size $b>2$ is useless as it increases the computational load while effectively reducing the amount of rotations being performed. Using blocks of size $2$ also allows us to use the explicit irreps representation of $2\times 2$ rotations and avoid computing a matrix exponential (eq.3). This also allows, as in RoPE su2023roformerenhancedtransformerrotary, to rewrite matrix multiplication of $R_{\theta}x$ in MapFormers as:
leaving only frequencies $\omega\in\mathbb{R}^{n_{b}}$ (see sec.10.2 for a discussion about their initialization) as learnable parameters instead of matrix $S\in\mathbb{R}^{n_{b}\times b\times b}$.
The actual rotation angle $\theta=\omega\Delta_{t}$ is dynamically computed from the input features $X$ using a low-rank projection $W_{\Delta}=W_{\Delta}^{out}W_{\Delta}^{in}$. The goal of this two-step process is to ensure the rotation relies only on the most salient action features of the input. First, the internal projection $W_{\Delta}^{in}\in\mathbb{R}^{d\times r}$ ($r\ll d$) maps the high-dimensional input $X$ to a low-dimensional $\Delta_{t}^{in}\in\mathbb{R}^{t\times r}$ (for instance, in a 2D environment, $\Delta_{t}^{in}$ could be the 2D movement vector $\Delta_{t_{k}}\in\mathbb{R}^{r}$ in eq. 3, where $r=2$). Second, this extracted action is projected by $W_{\Delta}^{out}$ into the final set of path increments $\Delta_{t}\in\mathbb{R}^{t\times n_{h}\times n_{b}}$, which controls the rotation angle for all attention heads ($n_{h}$) and all diagonal blocks ($n_{b}$), where $n_{b}=d_{h}/b$ is the number of blocks per head dimension.
Finally, compared to TEM and TEM-t that use a single neural population $p_{t}$ to encode position, our MapFormers use two separate initial vectors $k_{*}^{p}$ and $q_{*}^{p}$. This distinction is optional and mimics the formalism of EM-SSMs defined in eq. 10, meaning that we could set $k_{*}^{p}=q_{*}^{p}=p_{*}$ without loss of generality. However, we suspect this separation to be beneficial because it would create sparser attention values. Indeed, our positions $p_{t}$ vary smoothly through path-integration, implying that nearby positions would not be strongly separated, but introducing distinct neural populations for keys and queries could allow the model to find the optimal balance between separation/matching of distinct/similar positions.
5 Experiments
Figure 3: The 2D-Navigation Task and an Illustration of the Desired Corresponding Cognitive Map. (a) Illustration of the 2D navigation task: A model must predict the upcoming observation every time it comes back to it. The model only receives symbols, and must make sense of them without supervision. Some represent observations, others actions to take. (b) To solve this task, only actions (blue) should update the cognitive map, while observations (orange) leave it untouched. That is, observations should be projected to zero integration times. The logic is reversed for content, which is only updated by observation tokens, while actions are projected to zero (shaded blue). That is, value vectors of actions should be close to zero.
We defined different tasks that must all be solved without any supervision in order to demonstrate that with the appropriate neural architecture, prediction of upcoming information (next-token prediction) is enough for an agent to build a cognitive map. These tasks are sequences of symbols with a hidden structure that the model must infer in order to generalize.
To validate the computational benefits of rotation matrices (that will never be total, see sec. 10.4) we would like our models to allow the learning of lower level structures such as gating, while correctly representing 2-dimensional space. To test this versatility, we started with a 1-dimensional gating structure as our baseline task (section 5.1), as it is the most basic structure that a model should be able to represent, and since current selective models such as Mamba and CoPE manage to solve it, so should MapFormers. We then tested MapFormers on grid navigation, first in 1D, to show that even 1-dimensional structures are more than gating, and later in 2D, because it is the historical task Tolman1948-TOLCMI; O'Keefe1978; Hafting2005 and the most fundamental cognitive map to learn for an agent evolving in the physical world, and its irreducible representations are rotations that cannot be learned with the diagonal matrices of models listed above, which should demonstrate the superiority of MapFormers.
5.1 Selective copy
In this task introduced in sec.3.1, the model receives a sequence of symbols and must copy it entirely while ignoring a blank token $B$: $CFEABABBBCBF\to CFEAACF$, which requires the model to learn a gating mechanism to ignore blank tokens. OOD sequences are generated by varying the proportion of blank tokens. Our training sequences have 128/128 non blank/blank tokens, OOD Dense 128/64 and OOD sparse 128/256. For this task, we follow the training setup of golovneva2024contextualpositionencodinglearning and train models with 2 layers, 4 heads of size 64.
@@ 5.1 Selective copy
Table 1: Selective copy accuracy. IID sequences are generated with $128$ tokens to copy and $128$ blank tokens $B$ to ignore. OOD-d (dense) and OOD-s (sparse) vary the amount of blank tokens to ignore to $64$ and $256$, respectively. RoPE solves the task IID but fails to generalize OOD, TAPE improves generalization on shorter sequences but not on long sequences. As expected, both CoPE and MapFormer can solve it, showing that the latter can also learn gating mechanisms. MapEM-o and MapEM-s cannot solve this task, since it requires both "sensory" inputs (the tokens to copy) and structure (the tokens to ignore).
In line with golovneva2024contextualpositionencodinglearning, we find that RoPE perfectly learns in distribution but completely fails to generalize. TAPE outperforms RoPE on the OOD-dense dataset, implying that it has learned a form of position interpolation that RoPE cannot learn, but its performances collapse on the OOD-sparse dataset, implying that TAPE did not learn to integrate actions over long ranges. As expected, both CoPE and MapFormers manage to solve the task on OOD samples, since moving forward on a finite segment or a circle are equivalent. Because this task involves copying relevant tokens, it cannot be solved without sensory inputs (the token to copy), and therefore MapEM-s, our model relying on structure alone, cannot solve it.
This experiment shows that all methods learning cumulative actions (i.e. path integration, CoPE and MapFormers) manage to perfectly generalize to longer and shorter sequences.
5.2 Forced Grid Navigation
@@
Table 2: 1D-2D grid navigation accuracy. RoPE does not solve the task since it does not learn input-dependent matrices. CoPE performs better but does not solve the task, even in 1D, showing that gating is not enough to learn 1D cognitive maps, since it does not allow to learn the inverse of an action (forward/backward). Even with up to 4 layers, neither RoPE, TAPE nor CoPE manage to solve the task, showing that without the appropriate structural bias, deeper models cannot learn a cognitive map. All MapFormers model solved the task and perfectly generalize OOD, even if performances slightly degrade in 2D on longer sequences (OOD-s), since the amount of coordinates to represent increases quadratically. As expected MapEM-o, a transformer with no positional embeddings, performs even worse that RoPE, since it does not encode position.
An agent evolving in a two-dimensional world should understand how coordinates combine in order to keep track of its position. Our task consists of a random sequence of symbols $s=\left(a_{1},o_{1},a_{2},o_{2},...,a_{T},o_{T}\right)$ with action $a_{t}$ sampled from $\left\{\uparrow,\downarrow,\rightarrow,\leftarrow\right\}$ and observation $o_{t}$ from $\{o_{1},...,o_{K}\}$, a finite set of $K$ objects, plus a blank token $B$ to ignore. Every time the agent comes back to a previously visited location $p_{t}=p_{s},s<t\leq T$, it is asked to recall the last object $o_{s}$ it has seen there (fig. 3(a)).
In navigation tasks, we train all models on sequences of $l=128$ steps, $p_{empty}=0.5$ and a grid size of $l_{grid}=64$ and test length generalization on two OOD datasets - OOD-d: $l=64,p_{empty}=0.2,l_{grid}=32$ and OOD-s: $l=512,p_{empty}=0.8,l_{grid}=128$.
Because this task only requires recall of information, we hypothesize that one layer coupled with path integration should be enough, as in whittington2022relatingtransformersmodelsneural. We set the inner rank $r$ (see sec.4) of the integration time projection $W_{\Delta}$ to the world dimension $n$: $r:=n=1,2$ in 1D and 2D, respectively. We trained single layer models with two heads of size $h=64$ on 200K sequences, with AdamW optimizer loshchilov2019decoupledweightdecayregularization, a linear learning rate decay, a base learning rate $lr=3e^{-4}$, weight decay $0.05$ and a batch size of $128$ on a single GPU.
1D grid navigation
We first tested our models on a 1D navigation task, which has a subtle difference with gating. Even though the underlying structure is 1-dimensional, the model must be able to move forward or backward – i.e. representing an action and its inverse – while gating only compresses the ordinal order and can only move forward. Therefore, we hypothesize that CoPE, which learns a gating structure, should not be able to solve this task. As expected, all MapFormer models solve the task, but RoPE, CoPE and TAPE fail (tab. 2). The fact that CoPE fails in this task shows a key limitation of gating, even in 1D.
2D grid navigation
As expected, our three MapFormer variants perfectly solve the task in 2D and generalize to shorter or longer sequences, unlike other architectures. Even though CoPE does significantly better than RoPE, it fails to reach a perfect accuracy on both IID and OOD samples. Note that because 2D positions require more neurons to represent the two orthogonal axis, performances of MapEM and MapWM degrade faster on longer sequences than in 1D (tab. 2).
For pedagogical reasons, we added a control MapEM-o, a transformer with observation only, and as expected, the model does not learn anything, showing that structure is indeed learned through input-dependent positional encoding.
Increasing the number of layers for RoPE, CoPE, or TAPE does not change the model’s performance. CoPE seems to benefit from more layers in 1D only, probably because it helps the model to learn a backward operation that cannot be approximated with a single layer. TAPE’s performances improve IID with more layers, but increasing the model’s depth does not seem to impact OOD generalization. These results suggest that structure cannot be approximated with deeper networks without the appropriate inductive bias.
These experiments highlight the importance of learning input-dependent positional encoding in order to solve structure-dependent tasks with Transformers. CoPE, which only learns a gating mechanism, can solve the Selective Copy task but fails to learn a cognitive map, compared to MapFormers that perfectly generalize OOD in all tasks, demonstrating the expressive power of input-dependent matrices to solve structure-dependent problems.
@@
5.3 EM is computationally more expensive but scales better than WM
@@
Compared to MapFormer-WM, MapFormer-EM queries its memories with the flattened conjunction of observational and structural features, which requires a second pool of neurons $p_{t}\in\mathbb{R}^{d}$ as well as computing two attention matrices $A_{X}$ and $A_{P}$ separately. This offers sparser and higher dimensional features $g_{t}=\text{vec}(x_{t}.p_{t}^{\top})\in\mathbb{R}^{d^{2}}$ (fig. 1(c)) but doubles the computational load of the attention operation, which is already the main bottleneck in scaling transformers.
Figure 4: Behavioral Analyses: MapEM scales better than MapWM (a) Fixed sequence length $l=256$, varying head size. (b) Fixed head size $h=48$, varying sequence length. EM models are more robust to number of neurons and sequence length. (c) Fixed head size $h=32$ and sequence length $l=16$. Increasing the number of items to remember makes the task harder. Working memory models require more neurons to remember a large number of items.
This factorization in two separate pool of neurons should allow EM to be more efficient than WM, as in the former, neurons specialize for either position or observation. To verify our claim, we start by varying the number of neurons per head, to validate that MapEM models need less neurons, and in the same logic, we fix the model size but vary the sequence length to demonstrate that EM models represent positions with more precision. Finally, we decided to fix everything but vary the number of objects seen in training, to show that structure and memorization influence each other.
To test the impact of model size and sequence length, we vary the head size from $16$ to $128$ and sequence length from $16$ to $384$, respectively, while keeping a small vocabulary size of $10$. To test the impact of vocabulary on recall, we varied the number of objects from 10 to 10000, and train small models with head size 32 on short sequences of size 16. Results are reported in fig. 4.
@@
In line with Whittington2025, EM models scale better than WM, and unsurprisingly, our model relying on position alone (MapEM-s, blue) is more robust to both sequence length (fig. 4(a)) and model size (fig. 4(b)), since sensory features only add noise on this task. We clearly see here that the right architecture without enough neurons will completely fail to learn a cognitive map, highlighting the key role that model’s size plays in the ability to understand the structure of the world the agent evolves in.
The same effect is observed in memorization (fig. 4(c)), where WM models struggle to disambiguate objects when there are too many of them during training, while EM – which have a dedicated neural population for it – scale much better.
These experiments demonstrate the superior capacity of MapEM models in recall tasks, at the cost of doubling the attention computation.
5.4 Observations are vectors, actions are matrices
In our tasks, some symbols (actions $s_{t}=a_{t}\in\{\uparrow,\downarrow,\rightarrow,\leftarrow\}$) must update the model’s cognitive map (where), while others should leave it untouched (observations $s_{t}=o_{t}$), but modify its content (what). Since the choice of symbols is arbitrary, the model must learn by itself which ones are actions and which ones are observations. In order to demonstrate that the model has learned a cognitive map, we compare the learned representations of action versus observation tokens of our MapFormer models in the 2D Forced Navigation task. To demonstrate that the model has learned that only action tokens $a_{t}$ modify the cognitive map, we first compared the learned action representations $\Delta_{t}^{in}\in\mathbb{R}^{2}$ and $\Delta_{t}\in\mathbb{R}^{n_{h}\times n_{b}}$, since they define how the cognitive map must be updated. We also compared the representations of opposite actions (left/right and up/down), to verify that they cancel each other, as predicted by our theory. Finally, we also analyzed the learned value embeddings $v_{t}$ to demonstrate that only observation tokens change the state’s content.
@@ 5.4 Observations are vectors, actions are matrices
Therefore, in MapFormers, actions are matrices and observations are vectors, which resembles the ideas presented in baroni-zamparelli-2010-nouns, where nouns are vectors and adjectives are matrices, giving the model stronger OOD generalization compared to baseline models. Given the fact that Transformers are a generic architecture that can be used for language modeling, it is more than likely that the action/observation separation will focus on functional/content words, creating a separation between adjectives and verbs represented in the Lie algebra $\mathfrak{g}$ and nouns embedded in the content’s space.
Finally, the fact that observations $o_{t}$ do not update the cognitive map ($\Delta_{o_{t}}\approx 0$) and actions do not change its content ($v_{a_{t}}\approx 0$) implies that at each time step $t$, the model binds $o_{t}$ and $a_{t}$ together within the cognitive map and the content’s space, with position $p_{t}$ defined by $a_{t}$ and the updated value’s content $v^{\prime}_{t}$ by $o_{t}$. To give a better intuition, in MapEM-s – which uses position only to compute attention – attention updates value $v_{t}$ with a weighted sum of other values $v_{s\leq t}$ based on their structural similarity $A_{P}$. Since $o_{t}$ does not update the cognitive map, we have $p_{o_{t}}=R_{\theta_{o_{t}}}p_{a_{t}}\approx p_{a_{t}}:=p_{t}$, implying that their attention scores will be roughly equal. For the sake of simplicity, let us suppose that their similarity to other tokens in the sequence is null, meaning that $v^{\prime}_{a_{t}}=(v_{o_{t}}+v_{a_{t}})/2$ and $v^{\prime}_{o_{t}}=(v_{o_{t}}+v_{a_{t}})/2$. Because $v_{a_{t}}\approx 0$, we have $v^{\prime}_{o_{t}}=v^{\prime}_{a_{t}}=v_{o_{t}}/2:=v^{\prime}_{t}$. This means that MapFormers have the ability to flexibly group structurally similar tokens together (tokens $o_{t}$ and $a_{t}$ for all time steps $t$, fig.3(b)).
All these results are compelling evidence that our models have indeed learned to represent actions, and therefore a cognitive map.
@@ 6 Discussion
We introduced MapFormers, a novel class of Transformer-based architectures designed to learn cognitive maps and perform parallel path integration in a fully self-supervised setup. Our approach addresses the key limitation of previous RNN-based models by leveraging the parallel processing ability of Transformers, which is essential for scaling to complex, long-sequence problems. We developed two variants, MapFormer-WM (Working Memory) and MapFormer-EM (Episodic Memory), which unify relative and absolute positional embeddings, respectively, as computational models of these distinct memory systems. Critically, we provide theoretical arguments from Lie-group theory – extending the theoretical framework of actionable representations dorrell2023actionableneuralrepresentationsgrid – to understand the class of structures that will be represented within a MapFormer’s positional encoding (sec. 3.2 &\& 10.4)
Modern AI systems struggle with out-of-distribution (OOD) generalization and require orders of magnitude more data to achieve performance on par with humans, which can be explained by a misrepresentation of the true structure underlying the train and test data. Previous work tried to solve this issue in Transformers by learning input-dependent relative positional encoding (TAPE, CoPE and PoPE zhu2025rethinkingaddressinglanguagemodels; golovneva2024contextualpositionencodinglearning; gopalakrishnan2025decouplingwhatwherepolar), achieving robust length generalization on some structure dependent tasks, but provide no tool to understand which type of structures can or cannot be learned by the model. Moreover, CoPE and PoPE only adjust the relative distances between tokens, making them unsuitable for 2-dimensional tasks or beyond. TAPE cannot learn structure with a single layer since it updates its positional encoding after the attention, nor perform path integration by aggregating the cumulative effect of actions. Compared to these methods, our work: provides theoretical tools to understand what structure can or cannot be learned by the model; allows to represent actions explicitly in a dedicated neural population (the Lie algebra); does not focus uniquely on relative positional encoding, highlighting a core difference between absolute and relative positional encoding on recall tasks (sec. 5.3).
In our experiments, we found that MapFormers achieved almost perfect generalization to longer and sparser sequences, significantly outperforming baseline models. For problems that only required simple gating capabilities (sec. 5.1), MapFormers and CoPE golovneva2024contextualpositionencodinglearning achieved similar OOD generalization, as expected by our theory. TAPE zhu2025rethinkingaddressinglanguagemodels, which does not explicitly aggregate the effect of actions via a cumulative sum – i.e. does not learn to path integrate – fails to generalize OOD on all tasks, even with more layers. This highlights the importance of aggregating local effects via a cumulative sum to coherently represent structure, independently of sequence length. For more complex problems which require the creation of a cognitive map (sec.5.2), only MapFormers successfully generalized OOD. This demonstrates the important role of updating positional encoding with input-dependent matrices – endowed with the appropriate group structure – for structure-based OOD generalization, distinguishing it from generalization that might merely rely on sophisticated heuristics.
The key difference between EM and WM models comes from the explicit versus implicit representation of structure Whittington2025. In EM models, memories are formed by the conjunction of position and content, increasing the dimensionality of the representational space by a quadratic factor compared to WM models (as illustrated and reflected by the size of the final state in fig 1(c) &\&  1(d)). This specialization of neurons allows EM model to scale better, being more robust to model size, sequence length and vocabulary size (fig. 4).
How to know whether MapFormers have indeed learned a cognitive map? We identified several behavioral and neural evidences for that: First, the location in a cognitive map should be affected only by actions and not by content tokens (observations). fig. 5(a) shows that, indeed, once the model has learned that only actions should trigger rotations while observations should not, perfect generalization was achieved. Second, in a cognitive map, opposite actions (such as ‘step left’ vs. ‘step right’) should cancel each other. Indeed, fig. 5(b) and fig. 5(d) shows that the integration times (the coefficients of the representations of the actions in the Lie algebra) have opposite signs, and therefore would sum up to zero. Third, just as content tokens should not affect position in the cognitive map, actions should not have a ‘sensory’ content. Indeed, fig. 5(f) shows that the value vectors $V$’s of actions are an order-of-magnitude smaller than that of content observation tokens. Together, this shows that MapFormers learn to tell apart actions from observations, such that only actions affect the neural population that encode position in a cognitive map, with opposite effects for opposite actions, and only observations enter into content representation while actions have a negligible effect, allowing MapFormers to bind the position and content of structurally similar tokens. In other words, in MapFormers, observations are vectors and actions are matrices; MapFormers learn two distinct latent neural spaces: (1) a space that represents content (as in traditional deep learning models), and (2) a space that represents abstract actions in low dimension, within the Lie algebra (fig.1).
Compared to previous RNN-based EM/WM models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural, MapFormers exhibit several advantages. First, MapFormers overcome their computational bottleneck, requiring little extra compute compared to standard Transformers, maintaining their efficiency and scalability. Indeed, for problems that can be addressed with rotational operation (i.e., most cyclic abelian structures), using $2\times 2$ rotation matrices allowed us to avoid computing a matrix exponential and matrix-vector products (eq. 18), while remaining compatible with FlashAttention for MapWM models. Our experiments show the effectiveness of MapFormers, trained on significantly longer sequences than those explored in EM/WM RNNs so far. Furthermore, MapFormers provide greater flexibility compared to these models because they reduce the need to assume the underlying structure, and learn by themselves when to update or not the cognitive map. For instance, prior models Whittington2025; Whittington770495 required to explicitly know the number of possible actions and specify when to apply them, which is prohibitive in practice.
Second, the scalability and structural flexibility makes MapFormers, like Transformers, generic architectures that could be used in a variety of different tasks from language modeling to vision and protein folding siméoni2025dinov3; Jumper2021. The mechanism for learning a cognitive map – as our formalism shows – is also generic, which suggests that non-spatial cognitive maps can thus be learned and studied in a variety of different tasks. For example, in the case of language modeling, learning a cognitive map would mean disentangling the structure of sentences (syntax) from their content (semantics).
@@
Third, while we mainly explored rotation matrices for computational efficiency, but more complex group structures can be learned by MapFormers. This allows notably to explore structures where the irreducible representations (irreps dorrell2023actionableneuralrepresentationsgrid, sec. 3.2) are unknown, like rotations in high dimensions or non-commutative groups. This might affect inference speed; that is, a tradeoff between expressibility and computing time would need to be found (sec. 10.4).
Finally, MapFormers provide a biologically-plausible interpretation of their dynamics. The exponential map – needed to learn input-dependent matrices and therefore cognitive maps – arises in Mamba SSMs gu2024mambalineartimesequencemodeling; dao2024transformersssmsgeneralizedmodels and MapFormers via the integration of infinitesimal recurrent changes over a varying time interval. This explains how a neural population with fixed synaptic connections $A$ can learn a cognitive map by just varying the duration at which it is exposed to the input ($\overline{A_{t}}:=\exp(\Delta_{t}A)$; eq. 8). This suggests a biologically plausible neural mechanisms that allows animals to build cognitive maps, which could naturally emerge in continuous, selective recurrent networks.
@@
In summary, with MapFormers, we (1) address the issues of previous actionable models Whittington2025; Whittington770495; whittington2022relatingtransformersmodelsneural by enabling the learning of cognitive maps at scale; (2) unify absolute and relative positional encoding with models of Episodic Memory (EM) and Working Memory (WM); (3) give a comprehensive view of how to learn structure in Transformers, unlike previous attempts zhu2025rethinkingaddressinglanguagemodels; golovneva2024contextualpositionencodinglearning; gopalakrishnan2025decouplingwhatwherepolar; (4) empirically demonstrate that MapFormers are the only transformer-based architectures to perfectly generalize OOD on structure-dependent tasks.

### v3

@@ 4.2 Baseline models
In all experiments, we compare MapFormers to a classic Transformer architecture with static positional encoding, as well as two architectures learning input-dependent positional encoding serving as baselines for structure learning models.
@@
Rotary Position Embedding (RoPE) (Su et al., 2023) is used as our baseline Transformer since it is the most popular (fixed) position encoding architecture, and is analogous to our MapWM model without input selectivity, theoretically incapable of learning input-dependent matrices and structure-based problems.
Contextualized Positional Encoding (CoPE) (Golovneva et al., 2024) compresses relative distances with a cumulative gating mechanism, reminding path integration. Since the sigmoid gate can only take positive values, the model cannot represent an action and its inverse. Moreover, since CoPE only compresses relative ordinal distance, it should not be able to represent structure in higher dimensions.
@@
Contextualized Equivariant Position Embedding (TAPE) (Zhu et al., 2025) learns to modify its positional encoding after the attention mechanism, with input-dependent MLPs. However, TAPE exhibits several issues: First, updating structure after the attention means that the learned structure depends on the previous and not the current layer. Second, TAPE’s positional embedding are updated based on the attention scores and not path integration, giving no theoretical guaranty to integrate structure over long ranges.
MapEM-s and MapEM-o can choose to balance between observation $o$ and structure $s$, giving rise to MapEM-s, MapEM-o or MapEM-os by choosing to use either structure, observation or both in eq. 2, respectively.
5 Results
@@
Table 1: Selective copy accuracy. IID sequences have 128/128 blank/non-blank tokens. OOD: dense 64/128, sparse 256/128. RoPE solves the task IID but fails to generalize OOD, TAPE improves generalization on shorter sequences but not on long sequences. As expected, both CoPE and MapFormer can solve it, showing that the latter can also learn gating mechanisms. MapEM-o and MapEM-s cannot solve this task, since it requires both ”sensory” inputs (the tokens to copy) and structure (the tokens to ignore).
@@ D S D S
Table 2: 1D-2D grid navigation accuracy. IID sequences have 128/64 sequence length/grid width. OOD: dense (D) 64/32, sparse (S) 256/128. Neither RoPE and TAPE solve the task. CoPE performs better but also fails, even in 1D, since it cannot learn the inverse of an action (forward/backward). Even with additional layers compared to Mapformers, all baseline models fails. In contrast, MapFormers solved the task and perfectly generalized OOD, even if performances slightly degrade in 2D on longer sequences (OOD-s). This is since the amount of coordinates to represent increases quadratically with length. As expected, the control model MapEM-o performs even worse than RoPE, since it does not encode position.
5.1 MapFormers Solve the Selective-Copy Task
We start by evaluating MapFormers on the selective-copy task defined in sec. 3.1, compared to baseline models, following the training setup of (Golovneva et al., 2024). Since CoPE is able to solve this task by compressing token positions, we hypothesize that MapFormers will achieve similar results by encoding position on a 1D circle, using learned rotation. TAPE does not have theoretical guaranties for performing path integration, so we anticipated it would fail on OOD generalization. Tab. 1 shows that, indeed, only MapFormers and CoPE generalize OOD. Interestingly, TAPE improves performances in shorter sequences over RoPE, by learning a form of position interpolation, but performances deteriorate on longer sequences. In sum, these results show the importance of path integration for solving the selective-copy task.
5.2 MapFormers Solve the Grid Navigation Tasks but not Baseline Models
We next evaluated MapFormers on the navigation tasks in 1D and 2D (sec. 3.1). Because observations are random, to solve this task and generalize OOD a model needs to build coherent representations of actions and aggregate them sequentially. We therefore hypothesized that only MapFormers would achieve strong OOD generalization on these tasks given their ability to perform both path integration and motion on a torus, approximating a 2D plane. Since CoPE uses gating and cannot represent an action and its inverse, we expect it to fail even in 1D. Tab. 2 confirms that 1-layer MapFormers solve the task and generalize OOD, unlike baseline models, even with up to 4 layers. These results highlight the benefits of having a structural bias for path integration, and thus the ability to create a cognitive map of the external environment.
@@ 5.2 MapFormers Solve the Grid Navigation Tasks but not Baseline Models
5.3 MapEMs are more Computationally Demanding but scale better than MapWMs
@@
Compared to MapFormer-WM, MapFormer-EM queries its memories with the conjunction of observational and structural features, which requires a second pool of neurons $p_{t}\in\mathbb{R}^{d}$ as well as computing two attention matrices $A_{X}$ and $A_{P}$ separately. This offers sparser and higher dimensional features $g_{t}=\text{vec}(x_{t}.p_{t}^{\top})\in\mathbb{R}^{d^{2}}$ (fig. 5(a)) but doubles the computational load of the attention operation, which is already the main bottleneck in scaling transformers.
This factorization in two separate pools of neurons should allow EM to be more efficient than WM, as in the former, neurons specialize for either position or observation. To test this, we conducted three experiments to understand the effect of sequence length, model size and vocabulary size on performance, by varying each time one parameter while keeping the others fixed. Fig. 3 shows that, indeed, EM models scale better than WM, consistently with previous results (Whittington et al., 2025). Also, MapEM-s (blue curves), which relies on position alone, is more robust to both sequence length (fig. 3(a)) and model size (fig. 3(b)), since sensory features only add noise in this task. Last, having a sufficiently large number of neurons is crucial for both learning a cognitive map (fig. 3(a), models fail to converge with $h=16$) and memorizing large sets of objects (fig. 3(c)). In sum, these results demonstrate the superior capacity of MapEM models in recall tasks, at the cost of doubling the attention computation.
Figure 3: Behavioral Analyses: MapEM scales better than MapWM. Comparison of MapEM-s, MapEM-os, and MapWM robustness across model size, sequence length, and vocabulary size. (a) Train accuracy vs. head size (fixed sequence length $l=256$, vocabulary size $v=16$). MapEM requires fewer units than MapWM to achieve high performance. (b) Train accuracy vs. sequence length (fixed head size $h=48$, vocabulary $v=16$). EM models exhibit greater robustness to increasing sequence lengths. (c) Train accuracy vs. vocabulary size (fixed head size $h=32$, sequence length $l=16$). EM models outperform WM models in retaining an increasing number of items.
@@
5.4 Interpretability Analysis: Observations are Vectors, Actions are Matrices
In the tasks, some symbols (actions $s_{t}\in\{\uparrow,\downarrow,\rightarrow,\leftarrow\}$) must update the model’s position on its cognitive map, while others should leave position the same (observations) and instead modify its content. The model must learn by itself which ones are actions and which ones are observations. How can we tell whether MapFormers have learned this and therefore have learned a cognitive map?
@@ 5.4 Interpretability Analysis: Observations are Vectors, Actions are Matrices
From a point of view of Lie-group theory, actions are represented within the space of the Lie algebra $\mathfrak{g}$, whereas observations $v_{t}$ belong to the content space, the traditional latent space of transformers. Therefore, in MapFormers, actions are matrices and observations are vectors, which is reminiscent of ideas presented by Baroni and Zamparelli (2010), where nouns are vectors and adjectives are matrices, giving the model stronger OOD generalization compared to baseline models.
@@ 6 Discussion
We introduced MapFormers, a novel class of Transformer-based architectures designed to learn cognitive maps and perform parallel path integration in a self-supervised setup. MapFormers address the key limitations of previous RNN-based models in order to scale to complex, long-sequence problems: (1) they leverage the parallel processing ability of Transformers; (2) they offer greater flexibility by reducing the need to assume the underlying structure and learn when to update or not the cognitive map. We developed two variants, MapFormer-WM (Working Memory) and MapFormer-EM (Episodic Memory), which unify relative and absolute positional embeddings, respectively, as computational models of these distinct memory systems. Critically, we provide theoretical arguments from Lie-group theory—extending the theoretical framework of actionable representations (Dorrell et al., 2023)—to understand the class of structures that will be represented within a MapFormer’s positional encoding (sec. A.1 &\& B.2).
MapFormers achieved near-perfect OOD generalization on longer, sparser sequences, significantly outperforming baselines. While CoPE matched MapFormers on simple gating tasks – consistent with theoretical expectation – TAPE failed to generalize across all tasks since this type of model does not explicitly path integrate (sec. 5.1). Crucially, on complex tasks that require forming a cognitive map, only MapFormers succeeded (sec.5.2). This shows that updating positional encoding with input-dependent matrices—endowed with the appropriate group structure—is essential for true structural generalization rather than reliance on heuristics. Finally, probing the internal representations of the model, we found consistent evidence that MapFormers have indeed learned to represent a cognitive map.
@@
While we mainly explored rotation matrices for computational efficiency, more complex group structures can be learned by MapFormers. This allows notably to explore structures where the irreducible representations (irreps, Dorrell et al. 2023, sec. A.1) are unknown, like rotations in high dimensions or non-commutative groups. This might affect inference speed; that is, a tradeoff between expressibility and processing time needs to be found (sec. B.2).
@@
In summary, MapFormers (1) address the issues of previous actionable models (Whittington et al., 2019, 2022, 2025) by enabling the learning of cognitive maps at scale; (2) unify absolute and relative positional encoding with models of Episodic Memory (EM) and Working Memory (WM); (3) give a comprehensive view of how to learn structure in Transformers, unlike previous attempts (Zhu et al., 2025; Golovneva et al., 2024; Gopalakrishnan et al., 2025); (4) empirically demonstrate that MapFormers are the only current transformer-based architecture to perfectly generalize OOD on structure-dependent tasks.